OpenStudy (anonymous):

[4.03] Identify the correct slope and y-intercept of the equation y - 4x = -3 slope = -1; y-intercept = (0, -3) slope = -; y-intercept = (0, 3) slope = 4; y-intercept = (0, -3) slope = 4; y-intercept = (0, 3)

OpenStudy (eric_d):

y=mx+b

OpenStudy (eric_d):

b= y-intercept

OpenStudy (eric_d):

m=slope

OpenStudy (eric_d):

use this formula y=mx+b @kylieedeckahh

OpenStudy (bradely):

y - 4x = -3 Add 4x both sides y - 4x+4x = 4x-3 y=4x+(-3) The above equation in the form y=mx+c So m =4,c = -3 Option 3 is correct Source:www.mathskey.com

OpenStudy (eric_d):

Always remember y-intercept, x=0

OpenStudy (eric_d):

Got it @kylieedeckahh

OpenStudy (anonymous):

ohh that makes sense! thank you @eric_d

OpenStudy (eric_d):

yw